OpenStudy (anonymous):

If h(x) = -2x - 10, find h(-4). Answer -2 -3 -18 -16

OpenStudy (anonymous):

h(x) = -2x - 10, find h(-4). Substitute x=-4 h(-4)= -2(-4) - 10 = 8-10 =-2

OpenStudy (anonymous):

substitutee -4 into equation for x as tiapph has done